# Discrete Guidance Matching: Exact Guidance for Discrete Flow Matching

### v1

@@ 4.1 Simulation Results
Experimental Setup and Baselines. We consider a 2-D setting similar to that in zhang2025energy ; lu2023contrastive . The sample space is $\mathcal{S}^{\mathcal{D}}={\{0,1,\dots,32\}}^{2}$. We denote the source distribution as $p_{1}(x)$ and set the target distribution as its energy-guided version $p_{1}^{(\gamma)}(x)\propto p_{1}(x)e^{-\gamma\mathcal{E}(x)}$, where $\gamma\geq 0$ is the guidance strength and $\mathcal{E}(x)=-\log p(y=1|x)$ is the energy function defined by a given classifier $p(y=1|x)$. Our goal is to sample from the guided distribution $p_{1}^{(\gamma)}$ by using either the source transition rate $u^{p}_{t}$ or the source posterior $p_{1|t}$. We compare our proposed exact discrete guidance with the predictor guidance introduced in nisonoff2024unlocking . Note that the density ratio in this case is $r(x_{1})=\frac{p_{1}^{(\gamma)}(x_{1})}{p_{1}(x_{1})}=\frac{p^{\gamma}(y=1|x_{1})}{\mathcal{Z}(\gamma)}$, where $\mathcal{Z}(\gamma)$ is a constant that does not depend on $x_{1}$. Thus, we obtain the posterior-based and rate-based guidance based on Theorem 1 and Theorem 2, respectively. For the predictor guidance proposed by nisonoff2024unlocking , we use the predictor-guided rate $u_{t}^{(\gamma)}(z,x)={\Big[\frac{\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|z)}p(y=1|{\mathbf{x}}_{1})}{\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|x)}p(y=1|{\mathbf{x}}_{1})}\Big]}^{\gamma}u_{t}^{p}(z,x),~z\neq x$, which is different from our proposed rate-based guidance; see Section E.5 for further discussion.
Experimental Results. We sample from the 2-D target distribution $p_{1}^{(\gamma)}$ using different guidance schemes, as shown in Fig. 2 (a-c). For $\gamma=0,3$, both rate-based and posterior-based guidance produce distributions close to the ground truth. When $\gamma=10,20$, the distribution generated by the predictor guidance (nisonoff2024unlocking, ) differs from the ground-truth distribution. Overall, our posterior-based guidance achieves better performance than its rate-based counterpart, and also yields faster sampling, as illustrated in Fig. 2 (d). We include the experimental details in Section H.1 and more experimental results can be found in Appendix H.2.
@@
Figure 2: Comparison of selected sampling results (64 steps) in 2-D experiments and sampling efficiency. (a) Ground truth distribution; (b) rate-based guidance with masked initial distribution (nisonoff2024unlocking, ); (c) posterior-based guidance with uniform initial distribution (ours); (d) comparison of sampling times showing posterior-based guidance achieves 1.6x higher sampling speed than rate-based guidance.
@@ 4.2 RLHF on Multimodal Benchmark
RLHF Formulation. Let $\mathbf{c}$} denote the prompt with distribution $p_{\mathbf{c}}$, $\pi_{ref}({\mathbf{o}}_{1}|{\mathbf{c}})$ denote the reference policy associated with pretrained model $p^{d}_{1|t}({\mathbf{o}}_{1}^{d}|{\mathbf{c}},{\mathbf{o}}_{t})$. In the RL stage of RLHF (ouyang2022training, ), we consider to maximize
where $\tau$ is the temperature that controls the deviation from the reference policy. As mentioned in rafailov2023direct , the above objective has the closed-form maximizer
where $\mathcal{Z}({\mathbf{c}})=\int\pi_{ref}({\mathbf{o}}_{1}|{\mathbf{c}})\exp{\Big(\frac{\mathcal{R}({\mathbf{c}},{\mathbf{o}}_{1})}{\tau}\Big)}\text{d}{\mathbf{o}}_{1}$ and $\mathcal{R}(\cdot,\cdot)$ is the reward function which can be rule-based or obtained by training a neural network on the a comparison dataset. Given the reward function, we aim to generate samples ${\mathbf{o}}_{1}\sim\pi^{*}({\mathbf{o}}_{1}|{\mathbf{c}})$ conditional on $\mathbf{c}$}. If we take $p_{1}(\cdot)=\pi_{ref}(\cdot|{\mathbf{c}})$ and $q_{1}(\cdot)=\pi^{*}(\cdot|{\mathbf{c}})$ in Theorem 1, we can obtain a rectified posterior for sampling from $\pi^{*}$. We leave the detailed formulation in Appendix G.
Experimental Setup and Baselines. We develop the guidance based on a state-of-the-art discrete flow matching model on multimodal tasks (wang2025fudoki, ). For text-to-image generation tasks, we adopt the same training prompts and reward function as Liu2025FlowGRPOTF , which utilizes a weighted sum of Pickscore (kirstain2023pick, ) and GenEval reward (Ghosh2023GenEvalAO, ) as the reward function. We evaluate the performance on the widely used GenEval Benchmark (Ghosh2023GenEvalAO, ). For multimodal understanding tasks, we adopt the question and ground-truth answers from SEED (Li2023SEEDBenchBM, ) and the widely used LLM-as-a-judge for assigning reward. We adopt VLMEvalKit (Duan2024VLMEvalKitAO, ) to evaluate on several multimodal understanding datasets, including POPE (Li2023EvaluatingOH, ), MME-P (Fu2023MMEAC, ), MMB (Liu2023MMBenchIY, ), GQA (Hudson2019GQAAN, ), MMMU (Yue2023MMMUAM, ), and MM-Vet (Yu2023MMVetEL, ).
@@
Gen. Only LlamaGen (Sun2024AutoregressiveMB, ) 0.71 0.34 0.21 0.58 0.07 0.04 0.32
Emu3-Gen (Wang2024Emu3NP, ) 0.98 0.71 0.34 0.81 0.17 0.21 0.54
LDM (Rombach2021HighResolutionIS, ) 0.92 0.29 0.23 0.70 0.02 0.05 0.37
SDv1.5 (Rombach2021HighResolutionIS, ) 0.97 0.38 0.35 0.76 0.04 0.06 0.43
PixArt-alpha (Chen2023PixArtFT, ) 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SDv2.1 (Rombach2021HighResolutionIS, ) 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E 2 (Ramesh2022HierarchicalTI, ) 0.94 0.66 0.49 0.77 0.10 0.19 0.52
SDXL (Podell2023SDXLIL, ) 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E 3 (BetkerImprovingIG, ) 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium (Esser2024ScalingRF, ) 0.99 0.94 0.72 0.89 0.33 0.60 0.74
Und. + Gen. SEED- (Ge2024SEEDXMM, ) 0.97 0.58 0.26 0.80 0.19 0.14 0.49
LWM (Liu2024WorldMO, ) 0.93 0.41 0.46 0.79 0.09 0.15 0.47
ILLUME (Wang2024ILLUMEIY, ) 0.99 0.86 0.45 0.71 0.39 0.28 0.61
TokenFlow-XL (Qu2024TokenFlowUI, ) 0.95 0.60 0.41 0.81 0.16 0.24 0.55
Chameleon (Team2024ChameleonME, ) 0.39
Janus (Wu2024JanusDV, ) 0.97 0.68 0.30 0.84 0.46 0.42 0.61
Janus-Pro-1B (Chen2025JanusProUM, ) 0.98 0.82 0.51 0.89 0.65 0.56 0.73
Show-o (Xie2024ShowoOS, ) 0.95 0.52 0.49 0.82 0.11 0.28 0.53
Transfusion (Zhou2024TransfusionPT, ) 0.63
UniDisc (Swerdlow2025UnifiedMD, ) 0.92 0.47 0.15 0.67 0.13 0.19 0.42
D-DiT (Li2024DualDF, ) 0.97 0.80 0.54 0.76 0.32 0.50 0.65
FUDOKI (wang2025fudoki, ) 0.96 0.85 0.56 0.88 0.68 0.67 0.77
@@
Und. Only LLaVA-Phi-1.5 (Liu2023VisualIT, ) 1.3B 84.1 1128.0 - 56.5 30.7 -
MobileVLM (Chu2023MobileVLMA, ) 1.4B 84.5 1196.2 53.2 56.1 - -
MobileVLM-V2 (Chu2024MobileVLMVF, ) 1.4B 84.3 1302.8 57.7 59.3 - -
MobileVLM (Chu2023MobileVLMA, ) 2.7B 84.9 1288.9 59.6 59.0 - -
MobileVLM-V2 (Chu2024MobileVLMVF, ) 2.7B 84.7 1440.5 63.2 61.1 - -
LLaVA-Phi (Zhu2024LLaVAPhiEM, ) 2.7B 85.0 1335.1 59.8 - - 28.9
LLaVA (Liu2023VisualIT, ) 7B 76.3 809.6 38.7 - - 25.5
LLaVA-v1.5 (Liu2023ImprovedBW, ) 7B 85.9 1510.7 64.3 62.0 35.4 31.1
InstructBLIP (Dai2023InstructBLIPTG, ) 7B - - 36.0 49.2 - 26.2
Qwen-VL-Chat (Bai2023QwenVLAF, ) 7B - 1487.5 60.6 57.5 - -
IDEFICS (laurencon2023idefics, ) 8B - - 48.2 38.4 - -
Emu3-Chat (Wang2024Emu3NP, ) 8B 85.2 1244.0 58.5 60.3 31.6 37.2
InstructBLIP (Dai2023InstructBLIPTG, ) 13B 78.9 1212.8 - 49.5 - 25.6
Und. & Gen. LaVIT (Jin2023UnifiedLP, ) 7B - - - 46.8 - -
MetaMorph (Tong2024MetaMorphMU, ) 8B - - 75.2 - - -
Gemini-Nano-1 (Reid2024Gemini1U, ) 1.8B - - - - 26.3 -
ILLUME (Wang2024ILLUMEIY, ) 7B 88.5 1445.3 65.1 - 38.2 37.0
TokenFlow-XL (Qu2024TokenFlowUI, ) 13B 86.8 1545.9 68.9 62.7 38.7 40.7
LWM (Liu2024WorldMO, ) 7B 75.2 - - 44.8 - 9.6
VILA-U (Wu2024VILAUAU, ) 7B 85.8 1401.8 - 60.8 - 33.5
Chameleon (Team2024ChameleonME, ) 7B - - - - 22.4 8.3
Janus (Wu2024JanusDV, ) 1.5B 87.0 1338.0 69.4 59.1 30.5 34.3
Janus-Pro-1B (Chen2025JanusProUM, ) 1.5B 86.2 1444.0 75.5 59.3 36.3 39.8
Show-o-256 (Xie2024ShowoOS, ) 1.3B 73.8 948.4 - 48.7 25.1 -
Show-o-512 (Xie2024ShowoOS, ) 1.3B 80.0 1097.2 - 58.0 26.7 -
D-Dit (Li2024DualDF, ) 2.0B 84.0 1124.7 - 59.2 - -
FUDOKI (wang2025fudoki, ) 1.5B 86.1 1485.4 73.9 57.6 34.3 38.0
@@ 5 Related Works
Guidance. A line of work has studied guidance for diffusion and flow-matching models in continuous space (dhariwal2021diffusion, ; lu2023contrastive, ; ouyang2024transfer, ; feng2025guidance, ). Subsequent efforts extend guidance to discrete space (vignac2022digress, ; schiff2024simple, ; li2024derivative, ; nisonoff2024unlocking, ), but these methods typically rely on a first-order Taylor approximation to reduce the computational cost. However, this approximation is not appropriate in the discrete setting, which our work seeks to overcome.
Discrete Flow-based Models. Our work primarily builds on a line of discrete flow-based models (campbell2024generative, ; gat2024discrete, ; shaul2024flow, ; yimingdefog, ). These models construct transition rates that generate the target probability path through marginalizing conditional transition rate, yielding a more comprehensive design space of conditional probability path than discrete diffusion models (austin2021structured, ; campbell2022continuous, ; sun2022score, ; vignac2022digress, ; lou2023discrete, ) including masked diffusion models (shi2024simplified, ; sahoo2024simple, ; ou2024your, ; nie2025large, ; zhu2025llada, ; zhao2025d1, ; yang2025mmada, ). This motivates us to propose posterior-based guidance, which provides a flexible and unified perspective for generative modeling through marginalization.
@@ 6 Conclusion
In this work, we introduce a novel guidance framework for discrete flow-based models. Our approach provides exact guidance with improved sampling efficiency, offering a unified perspective for constructing guidance across generative models. We further propose training the guidance network by minimizing the Bregman divergence with regularization. The effectiveness of our framework is validated through experiments on energy-guided simulations, preference alignment for text-to-image generation, and multimodal understanding tasks.
@@ Appendix A Sampling Algorithms
Note that the transition probability (Equation 2) is valid only if $h\leq\frac{1}{{|u_{t}^{q}(x,x)|}}$ for each $x\in\mathcal{S}^{\mathcal{D}}$, when we just use its first-order approximation to generate samples. To remove this constraint, following the always-valid sampling procedure introduced in [65], given the current state ${\mathbf{x}}_{t}$, for each $d\in{[\mathcal{D}]}$, we first sample ${\mathbf{x}}_{1}^{d}$ from the learned posterior $q_{1|t}^{d,\theta}(\cdot|{\mathbf{x}}_{t})$, and then sample ${\mathbf{x}}_{t+h}^{d}$ from

### v2

@@ 4.1 Simulation Results
Experimental Setup and Baselines. We consider a 2-D setting similar to that in Zhang et al. (2025); Lu et al. (2023). The sample space is $\mathcal{S}^{\mathcal{D}}={\{0,1,\dots,32\}}^{2}$. We denote the source distribution as $p_{1}(x)$ and set the target distribution as its energy-guided version $p_{1}^{(\gamma)}(x)\propto p_{1}(x)e^{-\gamma\mathcal{E}(x)}$, where $\gamma\geq 0$ is the guidance strength and $\mathcal{E}(x)=-\log p(y=1|x)$ is the energy function defined by a given classifier $p(y=1|x)$. Our goal is to sample from the guided distribution $p_{1}^{(\gamma)}$ by using either the source transition rate $u^{p}_{t}$ or the source posterior $p_{1|t}$. We compare our proposed exact discrete guidance with the predictor guidance introduced in Nisonoff et al. (2025). Note that the density ratio in this case is $r(x_{1})=\frac{p_{1}^{(\gamma)}(x_{1})}{p_{1}(x_{1})}=\frac{p^{\gamma}(y=1|x_{1})}{\mathcal{Z}(\gamma)}$, where $\mathcal{Z}(\gamma)$ is a constant that does not depend on $x_{1}$. Thus, we obtain the posterior-based and rate-based guidance based on Theorem 1 and Theorem 2, respectively. For the predictor guidance proposed by Nisonoff et al. (2025), we use the predictor-guided rate $u_{t}^{(\gamma)}(z,x)={\Big[\frac{\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|z)}p(y=1|{\mathbf{x}}_{1})}{\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|x)}p(y=1|{\mathbf{x}}_{1})}\Big]}^{\gamma}u_{t}^{p}(z,x),~z\neq x$, which is different from our proposed rate-based guidance; see Section E.5 for further discussion.
Experimental Results. We sample from the 2-D target distribution $p_{1}^{(\gamma)}$ using different guidance schemes, as shown in Fig. 2 (a-c). For $\gamma=0,3$, both rate-based and posterior-based guidance produce distributions close to the ground truth. When $\gamma=10,20$, the distribution generated by the predictor guidance (Nisonoff et al., 2025) differs from the ground-truth distribution. Overall, our posterior-based guidance achieves better performance than its rate-based counterpart, and also yields faster sampling, as illustrated in Fig. 2 (d). We include the experimental details in Section H.1 and more experimental results can be found in Appendix H.3.
@@
Figure 2: Comparison of selected sampling results (64 steps) in 2-D experiments and sampling efficiency. (a) Ground truth distribution; (b) rate-based guidance with masked initial distribution (Nisonoff et al., 2025); (c) posterior-based guidance with uniform initial distribution (ours); (d) comparison of sampling times showing posterior-based guidance achieves 1.6x higher sampling speed than rate-based guidance.
@@ 4.2 RLHF on Multimodal Benchmark
Experimental Setup and Baselines. We develop the guidance based on a state-of-the-art discrete flow matching model on multimodal tasks, FUDOKI (Wang et al., 2025). The detailed probability path for FUDOKI can be found in equation 21 and the experimental details can be found in Section H.2. For text-to-image generation tasks, we adopt the same training prompts and reward function as Liu et al. (2025b), which utilizes a weighted sum of Pickscore (Kirstain et al., 2023) and GenEval reward (Ghosh et al., 2023) as the reward function. We evaluate the performance on the widely used GenEval Benchmark (Ghosh et al., 2023). For multimodal understanding tasks, we adopt the question and ground-truth answers from SEED (Li et al., 2023a) and the widely used LLM-as-a-judge for assigning reward. We adopt VLMEvalKit (Duan et al., 2024) to evaluate on several multimodal understanding datasets, including POPE (Li et al., 2023b), MME-P (Fu et al., 2023), MMB (Liu et al., 2023c), GQA (Hudson and Manning, 2019), MMMU (Yue et al., 2023), and MM-Vet (Yu et al., 2023).
@@
Gen. Only LlamaGen (Sun et al., 2024) 0.71 0.34 0.21 0.58 0.07 0.04 0.32
Emu3-Gen (Wang et al., 2024b) 0.98 0.71 0.34 0.81 0.17 0.21 0.54
LDM (Rombach et al., 2022) 0.92 0.29 0.23 0.70 0.02 0.05 0.37
SDv1.5 (Rombach et al., 2022) 0.97 0.38 0.35 0.76 0.04 0.06 0.43
PixArt-alpha (Chen et al., 2024) 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SDv2.1 (Rombach et al., 2022) 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E 2 (Ramesh et al., 2022) 0.94 0.66 0.49 0.77 0.10 0.19 0.52
SDXL (Podell et al., 2024) 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E 3 (Betker et al., ) 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium (Esser et al., 2024) 0.99 0.94 0.72 0.89 0.33 0.60 0.74
Und. + Gen. SEED- (Ge et al., 2024) 0.97 0.58 0.26 0.80 0.19 0.14 0.49
LWM (Liu et al., 2025a) 0.93 0.41 0.46 0.79 0.09 0.15 0.47
ILLUME (Wang et al., 2024a) 0.99 0.86 0.45 0.71 0.39 0.28 0.61
TokenFlow-XL (Qu et al., 2024) 0.95 0.60 0.41 0.81 0.16 0.24 0.55
Chameleon (Team et al., 2024) 0.39
Janus (Wu et al., 2025) 0.97 0.68 0.30 0.84 0.46 0.42 0.61
Janus-Pro-1B (Chen et al., 2025) 0.98 0.82 0.51 0.89 0.65 0.56 0.73
Show-o (Xie et al., 2025) 0.95 0.52 0.49 0.82 0.11 0.28 0.53
Transfusion (Zhou et al., 2025) 0.63
UniDisc (Swerdlow et al., 2025) 0.92 0.47 0.15 0.67 0.13 0.19 0.42
D-DiT (Li et al., 2024b) 0.97 0.80 0.54 0.76 0.32 0.50 0.65
FUDOKI (Wang et al., 2025) 0.96 0.85 0.56 0.88 0.68 0.67 0.77
@@
Und. Only LLaVA-Phi-1.5 (Liu et al., 2023b) 1.3B 84.1 1128.0 - 56.5 30.7 -
MobileVLM (Chu et al., 2023) 1.4B 84.5 1196.2 53.2 56.1 - -
MobileVLM-V2 (Chu et al., 2024) 1.4B 84.3 1302.8 57.7 59.3 - -
MobileVLM (Chu et al., 2023) 2.7B 84.9 1288.9 59.6 59.0 - -
MobileVLM-V2 (Chu et al., 2024) 2.7B 84.7 1440.5 63.2 61.1 - -
LLaVA-Phi (Zhu et al., 2024) 2.7B 85.0 1335.1 59.8 - - 28.9
LLaVA (Liu et al., 2023b) 7B 76.3 809.6 38.7 - - 25.5
LLaVA-v1.5 (Liu et al., 2023a) 7B 85.9 1510.7 64.3 62.0 35.4 31.1
InstructBLIP (Dai et al., 2023) 7B - - 36.0 49.2 - 26.2
Qwen-VL-Chat (Bai et al., 2023) 7B - 1487.5 60.6 57.5 - -
IDEFICS (Laurençon et al., 2023) 8B - - 48.2 38.4 - -
Emu3-Chat (Wang et al., 2024b) 8B 85.2 1244.0 58.5 60.3 31.6 37.2
InstructBLIP (Dai et al., 2023) 13B 78.9 1212.8 - 49.5 - 25.6
Und. & Gen. LaVIT (Jin et al., 2023) 7B - - - 46.8 - -
MetaMorph (Tong et al., 2024) 8B - - 75.2 - - -
Gemini-Nano-1 (et al., 2024) 1.8B - - - - 26.3 -
ILLUME (Wang et al., 2024a) 7B 88.5 1445.3 65.1 - 38.2 37.0
TokenFlow-XL (Qu et al., 2024) 13B 86.8 1545.9 68.9 62.7 38.7 40.7
LWM (Liu et al., 2025a) 7B 75.2 - - 44.8 - 9.6
VILA-U (Wu et al., 2024) 7B 85.8 1401.8 - 60.8 - 33.5
Chameleon (Team et al., 2024) 7B - - - - 22.4 8.3
Janus (Wu et al., 2025) 1.5B 87.0 1338.0 69.4 59.1 30.5 34.3
Janus-Pro-1B (Chen et al., 2025) 1.5B 86.2 1444.0 75.5 59.3 36.3 39.8
Show-o-256 (Xie et al., 2025) 1.3B 73.8 948.4 - 48.7 25.1 -
Show-o-512 (Xie et al., 2025) 1.3B 80.0 1097.2 - 58.0 26.7 -
D-Dit (Li et al., 2024b) 2.0B 84.0 1124.7 - 59.2 - -
FUDOKI (Wang et al., 2025) 1.5B 86.1 1485.4 73.9 57.6 34.3 38.0
@@ 5 Related Works
Guidance. A line of work has studied guidance for diffusion and flow-matching models in continuous space (Dhariwal and Nichol, 2021; Lu et al., 2023; Ouyang et al., 2024; Feng et al., 2025). Subsequent efforts extend guidance to discrete space (Vignac et al., 2023; Schiff et al., 2025; Li et al., 2024a; Nisonoff et al., 2025), but these methods typically rely on a first-order approximation to reduce the computational cost. However, this approximation is not appropriate in the discrete setting, which our work seeks to overcome.
Discrete Flow-based Models. Our work primarily builds on a line of discrete flow-based models (Campbell et al., 2024; Gat et al., 2024; Shaul et al., 2025; Qin et al., 2025). These models construct transition rates that generate the target probability path through marginalizing conditional transition rate, yielding a more comprehensive design space of conditional probability path than discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Sun et al., 2023; Vignac et al., 2023; Lou et al., 2024) including masked diffusion models (Shi et al., 2024; Sahoo et al., 2024; Ou et al., 2025; Nie et al., 2025b; Zhu et al., 2025; Zhao et al., 2025; Yang et al., 2025). This motivates us to propose posterior-based guidance, which provides a flexible and unified perspective for generative modeling through marginalization.
@@ 6 Conclusion
In this work, we introduce a novel guidance framework for discrete flow-based models. Our approach provides exact guidance with improved sampling efficiency, offering a unified perspective for constructing guidance across generative models. We further propose to train the guidance network by minimizing the Bregman divergence with regularization. The effectiveness of our framework is validated through experiments on energy-guided simulations, preference alignment for text-to-image generation, and multimodal understanding tasks.
Acknowledgments
Hongyuan Zha’s research is supported in part by Shenzhen Stability Science Program 2023, and National Natural Science Foundation of China (72495131). Fang Fang gratefully acknowledges research support from National Natural Science Foundation of China (72331005). Guang Cheng gratefully acknowledges financial support from the National Science Foundation (NSF–CNS 2247795).
Ethics statement
This work adheres to the ICLR Code of Ethics. Our research does not involve human subjects, sensitive personal data, or experiments that may pose harm to individuals or communities. Publicly available datasets and benchmarks were used in all experiments.
Reproducibility Statement
The code for our implementation can be found in the supplementary materials. For our theoretical result in Theorem 1 and Theorem 2, the assumption is stated in Assumption 1, and the proof can be found in Section D.2 and Section D.3.
@@ Appendix A Sampling Algorithms
Note that the transition probability (Equation 2) is valid only if $h\leq\frac{1}{{|u_{t}^{q}(x,x)|}}$ for each $x\in\mathcal{S}^{\mathcal{D}}$, when we just use its Euler discretization to generate samples. To remove this constraint, following the always-valid sampling procedure introduced in Shaul et al. (2025), given the current state ${\mathbf{x}}_{t}$, for each $d\in{[\mathcal{D}]}$, we first sample ${\mathbf{x}}_{1}^{d}$ from the learned posterior $q_{1|t}^{d,\theta}(\cdot|{\mathbf{x}}_{t})$, and then sample ${\mathbf{x}}_{t+h}^{d}$ from